OpenStudy (anonymous):

Find the dimensions of the rectangle if the width is 3 feet less than the length and the area is 40ft^2

OpenStudy (jdoe0001):

 \(\bf 40 = length \times (length - 3)\) solve for "length"

OpenStudy (anonymous):

(40) x (-3)?

myininaya (myininaya):

He is asking you to solve 40=l(l-3) for l You should recognize this is a quadratic equation. Try multiply l and (l-3) first I replace the length with l.

OpenStudy (anonymous):

Thats a one?

myininaya (myininaya):

A lowercase L. 1<--this is a one.

myininaya (myininaya):

40=L(L-3) You like L better? Either way, this is still a quadratic. Your objective is put into this form 0=ax^2+bx+c .

myininaya (myininaya):

or since we are using L instead of x into this form 0=aL^2+bL+c